# Quantum Kinetic Theory of the Linear Response for Weakly Disordered Multiband Systems

### v1

@@ Abstract
with $\theta_{\mu\nu}\equiv\theta(-\epsilon_{\mu}(\epsilon_{\mu}+\omega_{\nu}))$, by using the analytic structure of the Matsubara Green’s function [43]. From Eq.(7) with $n=m$, one obtains at the same leading order $f_{n}^{(1)}=-\partial_{\varepsilon}f_{n}^{(0)}\tau_{n}\sum_{i}(eE^{i})J^{i}_{n}$, which confirms the expected equivalence between Eq.(9) and Eq.(2), and provides an explicit expression for $J^{i}_{n}$ in terms of $f_{n}^{(1)}$. Then, going back to Eq.(7), it is clear that the interband components of the density matrix are either contributed by the interband components of the current vertex or by those of the two Green’s functions. Then, the first term on the RHS of Eq.(10) gives the first term on the RHS of Eq.(Quantum Kinetic Theory of the Linear Response for Weakly Disordered Multiband Systems). As for the third term on the RHS of Eq.(Quantum Kinetic Theory of the Linear Response for Weakly Disordered Multiband Systems), the second term of Eq.(10)) yields the terms proportional to $\underaccent{\tilde}{f}^{(1)}_{q}$, whereas the interband components of the Green’s functions, as derived from Eq.(8), lead to the terms proportional to $f^{(1)}_{n}$ and $f^{(1)}_{m}$ [43]. This completes the demonstration of the equivalence between our formalism and the Kubo formula in the ladder approximation, in the constant and homogeneous limits. One needs of course to add that our formalism also validly captures the part of the linear responses slowly varying in time and space at the same leading orders in the semiclassical and disorder expansions, in the case of finite mesoscopic systems. This is much more challenging to achieve with fully quantum calculations.
Pedagogical application to SOT in the 2DEG– As a pedagogical example, we address now the microscopic theory of SOT in the weakly disordered 2DEG, a topic previously investigated with the Kubo formula in the ladder approximation in [20]. The model Hamiltonian is
@@
with ${\bm{p}}\equiv(p^{x},p^{y})$, $\varepsilon_{0}$ some paramagnetic band dispersion, $J_{e}>0$ the exchange constant, ${\bm{m}}$ a classical unit vector aligned with the magnetization and ${\bm{b}}_{\scriptscriptstyle SO}$ some momentum dependent SO field with $b_{\scriptscriptstyle SO}\ll J_{e}$. The operator $\hat{\bm{\sigma}}$ is the vector of Pauli matrices, while $\hat{\sigma}_{0}$ is the identity matrix in spin space. The Hamiltonian (11) is a generalization, to arbitrary band and SO field dispersions, of the model considered in [20]. The dynamics of ${\bm{m}}$ is controlled by the Landau-Lifschitz-Gilbert equation [61], which in the dissipationless limit and spatially uniform case reads $(M_{\scriptscriptstyle S}/\gamma)\partial_{t}{\bm{m}}={\bm{m}}\times(\delta\mathcal{F}/\delta{\bm{m}})\equiv{\bm{T}}$, with $\gamma(<0)$ the gyromagnetic ratio, $M_{\scriptscriptstyle S}$ the saturation magnetization, $\mathcal{F}$ the thermodynamic potential and ${\bm{T}}$ the torque density. The SOT, in linear response to a quasi uniform and constant electric field, is then given by ${\bm{T}}^{(1)}={\bm{m}}\times(\delta\mathcal{F}^{(1)}/\delta{\bm{m}})$, with $\mathcal{F}^{(1)}$ the linear change in $\mathcal{F}$. By application of the Hellmann-Feynman theorem [62, 63] at finite temperature [64], we then derive ${\bm{T}}^{(1)}={\bm{m}}\times\langle\delta\hat{h}/\delta{\bm{m}}\rangle^{(1)}=J_{e}{\bm{m}}\times\langle\hat{\bm{\sigma}}\rangle^{(1)}$. The band energies and associated projectors for the Hamiltonian (11) are trivially obtained as $\varepsilon_{\sigma}=\varepsilon^{0}+\sigma\Delta$ and $\hat{P}_{\sigma}=\frac{1}{2}\left(\hat{\sigma}_{0}+\sigma{\bm{u}}\cdot\hat{\bm{\sigma}}\right)$, respectively ; in which we have introduced $\sigma\in\{+1,-1\}\equiv\{\uparrow,\downarrow\}$, with $\Delta\approx J_{e}(1+{\bm{m}}\cdot\underaccent{\bar}{\bm b}_{\scriptscriptstyle SO})$ and ${\bm{u}}\approx{\bm{m}}+{\bm{m}}\times(\underaccent{\bar}{\bm b}_{\scriptscriptstyle SO}\times{\bm{m}})$, as we are approximating from now on all quantities at leading order in $\underaccent{\bar}{\bm b}_{\scriptscriptstyle SO}={\bm{b}}_{\scriptscriptstyle SO}/J_{e}$. The interband Berry connection reads $\hat{\mathcal{A}}^{i}_{\sigma\underaccent{\tilde}{\sigma}}\ \approx\ \frac{{\bf i}\underaccent{\tilde}{\sigma}}{4}\partial_{p^{i}}\underaccent{\bar}{\bm b}_{\perp}\cdot\hat{\bm{\sigma}}\ +\ \frac{1}{4}({\bm{m}}\times\partial_{p^{i}}\underaccent{\bar}{\bm b}_{\perp})\cdot\hat{\bm{\sigma}}$, with $\underaccent{\bar}{\bm b}_{\perp}={\bm{m}}\times(\underaccent{\bar}{\bm b}_{\scriptscriptstyle SO}\times{\bm{m}})$. The intrinsic contribution to the SOT, from the first term on the RHS of Eq.(Quantum Kinetic Theory of the Linear Response for Weakly Disordered Multiband Systems), immediately ensues as
@@
${\bm{T}}^{(1)}_{int}\approx\frac{\hbar}{2}\int_{p}\left(f^{(0)}_{\uparrow}-f^{(0)}_{\downarrow}\right)(e{\bm{E}})\cdot\partial_{\bm{p}}\underaccent{\bar}{\bm b}_{\perp}.$ (12)
@@
${\bm{T}}^{(1)}_{int}\approx\frac{\hbar}{2}\sum_{\sigma}\sigma\int\displaylimits_{FS_{\sigma}}\!\!\frac{dl_{p}}{(2\pi\hbar)^{2}}\ \frac{(e{\bm{E}})\cdot\partial_{\bm{p}}\varepsilon_{0}}{|\partial_{\bm{p}}\varepsilon_{0}|}\underaccent{\bar}{\bm b}_{\perp},$ (13)
@@
with $dl_{p}$ being the infinitesimal line element along the 1D Fermi surface for band $\sigma$ in absence of SO interaction i.e., $FS_{\sigma}$ 111 In the case of a Fermi surface intersecting the Brillouin zone (BZ) boundary, the line integrals over some portions of this boundary resulting from the application of the divergence theorem does not contribute to the RHS of Eq.(12). Indeed, since the 2D BZ is by definition a primitive unit cell of the reciprocal space over which $\hat{h}$ is periodic, such contributions will cancel out by pairs, from pairs of line segments on opposite sides of the BZ boundary, one being the translation of the other by a primitive reciprocal lattice vector. Hence, we have established that under the stated hypothesis, the intrinsic SOT, which according to Eq.(12) would be classified as a Fermi sea contribution, can equally be expressed as a Fermi surface one according to Eq.(13). This exactly parallels the classical result obtained by Haldane in the case of the non quantized part of the AHE conductivity [44]. Having established this non trivial result, we focus now on the disorder correction to ${\bm{T}}^{(1)}_{int}$, which is stemming from the third term on the RHS of our Eq.(Quantum Kinetic Theory of the Linear Response for Weakly Disordered Multiband Systems). Specializing it to the case at hand, we readily obtain this contribution to the interband Wigner function as
@@
$\!\!\!\!\hat{\rho}^{(1)}_{dis}\!\approx\!\frac{\pi n_{i}v_{0}^{2}}{2J_{e}}{\bm{m}}\times\!\sum_{\sigma,\underaccent{\tilde}{\sigma}}\underaccent{\tilde}{\sigma}\!\!\!\!\int\displaylimits_{{FS}_{\scriptscriptstyle\underaccent{\tilde}{\sigma}}}\!\!\!\!\!\frac{dl_{\underaccent{\tilde}{p}}}{(2\pi\hbar)^{2}}\!\frac{f_{\sigma}^{\scriptscriptstyle(1)}\!-\!\underaccent{\tilde}{f}_{\underaccent{\tilde}{\sigma}}^{\scriptscriptstyle(1)}}{|\partial_{\underaccent{\tilde}{\bm p}}\underaccent{\tilde}{\varepsilon}_{0}|}(\underaccent{\bar}{\bm b}_{\perp}\!-\underaccent{\tilde}{\underaccent{\bar}{\bm b}}_{\perp})\cdot\hat{\bm{\sigma}},$ (14)
@@
with the distribution function to be derived without SO interaction. In this limit, the scattering Kernel $W_{\sigma,\underaccent{\tilde}{\sigma}}$ becomes diagonal and the BEs for the two spin bands decouple. At this point, we further specialize the problem to a parabolic 2DEG as in [20] i.e., to $\varepsilon_{0}\equiv p^{2}/(2m^{\star})$ with $m^{\star}$ the effective mass. One can then readily solves the BEs with $f_{\sigma}^{(1)}=\delta\left({\varepsilon}_{\sigma}-\varepsilon_{\scriptscriptstyle F}\right)(\tau_{F}/m^{\star})(e{\bf E})\cdot{\bm{p}}$, with the relaxation time and the density of states at the Fermi level respectively given by $1/\tau_{F}=(2\pi n_{i}v_{0}^{2}D_{F})/\hbar$ and $D_{F}=m^{\star}/(2\pi\hbar^{2})$. Substituting this into Eq.(14), and computing the associated spin density, leads to
@@
${\bm{T}}^{(1)}_{dis}\approx-\frac{\hbar}{2}\sum_{\sigma}\sigma\int\displaylimits_{FS_{\sigma}}\!\!\frac{dl_{p}}{(2\pi\hbar)^{2}}\ (e{\bm{E}})\cdot\frac{\bm{p}}{p}\ \underaccent{\bar}{\bm b}_{\perp},$ (15)
@@
Conclusions – We have provided a quantum kinetic formulation of the linear response to a an electric field slowly varying in time and space, for a general multiband electron system in presence of weak scalar disorder. While our formalism constitutes a theoretical and computational alternative to the Kubo formula in the stated context, with distinct advantages for finite systems and in capturing disorder effects beyond the CBA, we have proven the equivalence of the two formulations at the ladder approximation level in the homogeneous case. As a pedagogical example, we applied our approach to SOT theory for a simple two-band model of 2DEG amenable to analytical calculations, recovering and generalizing with striking efficiency a major vertex correction, while uncovering a new instance of intrinsic response as a Fermi surface property. We shall stress that our formalism can be applied with equal effectiveness to more involved multiband models up to density functional theory generated tight-binding Hamiltonian, however requiring in such case a numerical resolution of the Boltzmann equation controlling the intraband response. We are already undertaking such studies, for instance to shed new light on the influence of disorder on the SOT in van der Waals ferromagnets exhibiting anisotropic band structure with $C_{3}v$ symmetry [66], a subject of high current interest [18]. It shall also be noted that the disorder model can be easily extended through the consideration of proper self-energies, for instance to include SO scattering contributions. Finally, we want to mention some potential important implications of our formalism on another timely topic i.e., the theory of out-of-equilibrium orbital angular momentum (OAM) in time reversal and inversion symmetric metals, as a possible common explanation for several recently observed effects [47, 45, 46]. In this context, it is to be noted that most Kubo formula based studies have focused so far on the computation of the intrinsic OHE conductivity in the bulk [14, & references therein], irrespective of the chosen definition for the OAM observable. In our formulation, the intrinsic OHE current can be recovered from the first term on the RHS of Eq.(Quantum Kinetic Theory of the Linear Response for Weakly Disordered Multiband Systems). Beyond the possible fragility of this current under disorder corrections [21], which can be straightforwardly assessed from the third term on the RHS of our Eq.(Quantum Kinetic Theory of the Linear Response for Weakly Disordered Multiband Systems), we want to point out the importance of the second term i.e., of quantum coherences generated by the gradients in the distribution functions. It can be shown [67] that this is a source of orbital edge accumulation, that is only partially correlated to the OHE conductivity through the Berry connection and which may be an important explanatory factor to the previously cited experimental observations.

### v2

@@ Abstract
with $\theta_{\mu\nu}\equiv\theta(-\epsilon_{\mu}(\epsilon_{\mu}+\omega_{\nu}))$, by using the analytic structure of the Matsubara Green’s function [46]. From Eq.(7) with $n=m$, one obtains at the same leading order $f_{n}^{\scriptscriptstyle(1)}=-\partial_{\varepsilon}f_{n}^{\scriptscriptstyle(eq)}\tau_{n}\sum_{i}(eE^{i})J^{i}_{n}$, which confirms the expected equivalence between Eq.(9) and Eq.(2), and provides an explicit expression for $J^{i}_{n}$ in terms of $f_{n}^{\scriptscriptstyle(1)}$. Then, going back to Eq.(7), it is clear that the interband components of the density matrix are either contributed by the interband components of the current vertex or by those of the two Green’s functions. Then, the first term on the RHS of Eq.(10) gives the first term on the RHS of Eq.(Quantum Kinetic Theory of the Linear Response for Weakly Disordered Multiband Systems). As for the third term on the RHS of Eq.(Quantum Kinetic Theory of the Linear Response for Weakly Disordered Multiband Systems), the second term of Eq.(10)) yields the terms proportional to $\underaccent{\tilde}{f}^{\scriptscriptstyle(1)}_{q}$, whereas the interband components of the Green’s functions, as derived from Eq.(8), lead to the terms proportional to $f^{\scriptscriptstyle(1)}_{n}$ and $f^{\scriptscriptstyle(1)}_{m}$ [46]. This completes the demonstration of the equivalence between our formalism and the Kubo formula in the ladder approximation, in the constant and homogeneous limits. One needs of course to add that our formalism also validly captures the part of the linear responses slowly varying in time and space at the same leading orders in the semiclassical and disorder expansions, in the case of finite mesoscopic systems. This is much more challenging to achieve with fully quantum calculations.
Pedagogical application to the SOT in the 2DEG– As a pedagogical example, we address now the microscopic theory of the SOT in the weakly disordered 2DEG, a topic previously investigated with the Kubo formula [20, 65]. The model Hamiltonian is
@@
with ${\bm{p}}\equiv(p^{x},p^{y})$, $\varepsilon_{0}$ some paramagnetic band dispersion, $J_{e}>0$ the exchange constant, ${\bm{m}}$ a classical unit vector aligned with the magnetization and ${\bm{b}}_{\scriptscriptstyle SO}$ some momentum dependent SO field with $b_{\scriptscriptstyle SO}\ll J_{e}$. The operator $\hat{\bm{\sigma}}$ is the vector of Pauli matrices, while $\hat{\sigma}_{0}$ is the identity matrix in spin space. The Hamiltonian (11) is a generalization, to arbitrary band and SO field dispersions, of the model considered in [20]. If we neglect internal dissipation [66, 67], the dynamics of ${\bm{m}}$ is controlled by the Landau-Lifschitz equation [68, 69] i.e., $(M_{\scriptscriptstyle S}/\gamma)\partial_{t}{\bm{m}}={\bm{m}}\times(\delta\mathcal{F}/\delta{\bm{m}})\equiv{\bm{T}}$, with $\gamma(<0)$ the gyromagnetic ratio, $M_{\scriptscriptstyle S}$ the saturation magnetization, $\mathcal{F}$ the thermodynamic potential and ${\bm{T}}$ the torque density. The SOT, in linear response to a quasi uniform and constant electric field, is then given by ${\bm{T}}^{(1)}={\bm{m}}\times(\delta\mathcal{F}^{\scriptscriptstyle(1)}/\delta{\bm{m}})$, with $\mathcal{F}^{\scriptscriptstyle(1)}$ the linear change in $\mathcal{F}$. By application of the Hellmann-Feynman theorem [70, 71] at finite temperature [72], and assuming a spatially uniform magnetizartion, we then obtain ${\bm{T}}^{\scriptscriptstyle(1)}={\bm{m}}\times\langle\delta\hat{h}/\delta{\bm{m}}\rangle^{\scriptscriptstyle(1)}=J_{e}{\bm{m}}\times\langle\hat{\bm{\sigma}}\rangle^{\scriptscriptstyle(1)}$. The band energies and associated projectors for the Hamiltonian (11) are trivially obtained as $\varepsilon_{\sigma}=\varepsilon^{0}+\sigma\Delta$ and $\hat{P}_{\sigma}=\frac{1}{2}\left(\hat{\sigma}_{0}+\sigma{\bm{u}}\cdot\hat{\bm{\sigma}}\right)$, respectively ; in which we have introduced $\sigma\in\{+1,-1\}\equiv\{\uparrow,\downarrow\}$, with $\Delta\approx J_{e}(1+{\bm{m}}\cdot\underaccent{\bar}{\bm b}_{\scriptscriptstyle SO})$ and ${\bm{u}}\approx{\bm{m}}+{\bm{m}}\times(\underaccent{\bar}{\bm b}_{\scriptscriptstyle SO}\times{\bm{m}})$, as we are approximating from now on all quantities at leading order in $\underaccent{\bar}{\bm b}_{\scriptscriptstyle SO}={\bm{b}}_{\scriptscriptstyle SO}/J_{e}$. The interband Berry connection reads $\hat{\mathcal{A}}^{i}_{\sigma\underaccent{\tilde}{\sigma}}\ \approx\ \frac{{\bf i}\underaccent{\tilde}{\sigma}}{4}\partial_{p^{i}}\underaccent{\bar}{\bm b}_{\perp}\cdot\hat{\bm{\sigma}}\ +\ \frac{1}{4}({\bm{m}}\times\partial_{p^{i}}\underaccent{\bar}{\bm b}_{\perp})\cdot\hat{\bm{\sigma}}$, with $\underaccent{\bar}{\bm b}_{\perp}={\bm{m}}\times(\underaccent{\bar}{\bm b}_{\scriptscriptstyle SO}\times{\bm{m}})$. The intrinsic contribution to the SOT, from the first term on the RHS of Eq.(Quantum Kinetic Theory of the Linear Response for Weakly Disordered Multiband Systems), immediately ensues as
@@
${\bm{T}}^{\scriptscriptstyle(1)}_{int}\approx\frac{\hbar}{2}\int_{p}\left(f^{\scriptscriptstyle(eq)}_{\uparrow}-f^{\scriptscriptstyle(eq)}_{\downarrow}\right)(e{\bm{E}})\cdot\partial_{\bm{p}}\underaccent{\bar}{\bm b}_{\perp}.$ (12)
@@
${\bm{T}}^{\scriptscriptstyle(1)}_{int}\approx\frac{\hbar}{2}\sum_{\sigma}\sigma\int\displaylimits_{FS_{\sigma}}\!\!\frac{dl_{p}}{(2\pi\hbar)^{2}}\ \frac{(e{\bm{E}})\cdot\partial_{\bm{p}}\varepsilon_{0}}{|\partial_{\bm{p}}\varepsilon_{0}|}\underaccent{\bar}{\bm b}_{\perp},$ (13)
@@
with $dl_{p}$ being the infinitesimal line element along the 1D Fermi surface for band $\sigma$ in the absence of SO interaction i.e., $FS_{\sigma}$ 222 If the Fermi surface intersects the Brillouin zone (BZ) boundary, the line integrals over some portions of this boundary resulting from the application of the divergence theorem does not contribute to the RHS of Eq.(12). Indeed, since the 2D BZ is by definition a primitive unit cell of the reciprocal space over which $\hat{h}$ is periodic, such contributions will cancel out by pairs, from pairs of line segments on opposite sides of the BZ boundary, one being the translation of the other by a primitive reciprocal lattice vector. Hence, we have established that under the stated hypothesis, the intrinsic SOT, which according to Eq.(12) would be classified as a Fermi sea contribution, can equally be expressed as a Fermi surface one according to Eq.(13). This exactly parallels the classical result obtained by Haldane in the case of the non quantized part of the AHE conductivity [47]. Having established this non trivial result, we focus now on the disorder correction to ${\bm{T}}^{\scriptscriptstyle(1)}_{int}$, which is stemming from the third term on the RHS of our Eq.(Quantum Kinetic Theory of the Linear Response for Weakly Disordered Multiband Systems). Specializing it to the case at hand, we readily obtain this contribution to the interband Wigner function as
@@
$\!\!\!\!\hat{\rho}^{\scriptscriptstyle(1)}_{dis}\!\approx\!\frac{\pi n_{i}v_{0}^{2}}{2J_{e}}{\bm{m}}\times\!\sum_{\sigma,\underaccent{\tilde}{\sigma}}\underaccent{\tilde}{\sigma}\!\!\!\!\int\displaylimits_{{FS}_{\scriptscriptstyle\underaccent{\tilde}{\sigma}}}\!\!\!\!\!\frac{dl_{\underaccent{\tilde}{p}}}{(2\pi\hbar)^{2}}\!\frac{f_{\sigma}^{\scriptscriptstyle(1)}\!-\!\underaccent{\tilde}{f}_{\underaccent{\tilde}{\sigma}}^{\scriptscriptstyle(1)}}{|\partial_{\underaccent{\tilde}{\bm p}}\underaccent{\tilde}{\varepsilon}_{0}|}(\underaccent{\bar}{\bm b}_{\perp}\!-\underaccent{\tilde}{\underaccent{\bar}{\bm b}}_{\perp})\cdot\hat{\bm{\sigma}},$ (14)
@@
with the distribution function to be derived without SO interaction. In this limit, the scattering Kernel $W_{\sigma,\underaccent{\tilde}{\sigma}}$ becomes diagonal and the BEs for the two spin bands decouple. At this point, we further specialize the problem to a parabolic 2DEG as in [20] i.e., to $\varepsilon_{0}\equiv p^{2}/(2m^{\star})$ with $m^{\star}$ the effective mass. One can then readily solve the BEs with $f_{\sigma}^{\scriptscriptstyle(1)}=\delta\left({\varepsilon}_{\sigma}-\varepsilon_{\scriptscriptstyle F}\right)(\tau_{F}/m^{\star})(e{\bf E})\cdot{\bm{p}}$, with the relaxation time and the density of states at the Fermi level respectively given by $1/\tau_{F}=(2\pi n_{i}v_{0}^{2}D_{F})/\hbar$ and $D_{F}=m^{\star}/(2\pi\hbar^{2})$. Substituting this into Eq.(14), and computing the associated spin density, leads to
@@
${\bm{T}}^{\scriptscriptstyle(1)}_{dis}\approx-\frac{\hbar}{2}\sum_{\sigma}\sigma\int\displaylimits_{FS_{\sigma}}\!\!\frac{dl_{p}}{(2\pi\hbar)^{2}}\ (e{\bm{E}})\cdot\frac{\bm{p}}{p}\ \underaccent{\bar}{\bm b}_{\perp},$ (15)
@@
Implications for the theory of orbital effects – Regarding the theory of out-of-equilibrium orbital angular momentum (OAM) generation in time reversal and inversion symmetric metals, as an explanation for several recently observed effects [50, 48, 49], most Kubo formula based studies have focused so far on the computation of the intrinsic OHE conductivity in the bulk [14, & references therein], irrespective of the chosen definition for the OAM observable. In our formulation, the intrinsic OHE current can be recovered from the first term on the RHS of Eq.(Quantum Kinetic Theory of the Linear Response for Weakly Disordered Multiband Systems). Beyond the possible fragility of this current under disorder corrections [21], which can be straightforwardly assessed from the third term on the RHS of our Eq.(Quantum Kinetic Theory of the Linear Response for Weakly Disordered Multiband Systems), we want to point out again the importance of the second term i.e., of quantum coherences generated by the gradients. It can be shown [74] that this is a source of orbital edge accumulation, that is only partially correlated to the OHE conductivity through the Berry connection and which may be an important explanatory factor to the previously cited experimental observations.
Conclusions – We provide a quantum kinetic theory of the linear response to an electric field slowly varying in time and space, for multiband electron systems in presence of weak scalar disorder. We uncovered gradients in the intraband response on mesoscopic scales as a source of interband coherence, opening a new research area as it suggests alternate ways to locally modulate current induced responses e.g., spin or orbital accumulation, by controlling the boundaries shape or scattering properties. While our theory constitutes an alternative to the Kubo formula in the stated context, with distinct advantages for finite systems and in going beyond the CBA, we have proven the equivalence of the two formulations at the ladder approximation level in the homogeneous case. We applied our approach to SOT theory for a simple model of 2DEG, recovering and generalizing with striking efficiency a major vertex correction, while uncovering a new instance of intrinsic response as a Fermi surface property. We also suggest possible implications for the theory of orbital effects. Our formalism is applicable to realistic material models, up to ab-initio generated tight-binding Hamiltonian, through numerical resolution of the Boltzmann equation. We are already undertaking such studies in relation to SOT in van der Waals magnets with $C_{3}v$ symmetry [75], a subject of high current interest [18]. Finally, the disorder model can be easily extended, for instance to include SO scattering contributions, and we also intend to address sub-leading contributions to the intraband response in a future work.